2) An economist believes there is a linear relationship between the market price p of a particular
commodity and the number of suppliers are willing to bring to the market place. Two sample
observations indicate that when the price equals $15 per unit, the weekly supply equals 30,000
units and when the price is $20 per unit the weekly supply is 48000 units.
i) If price p is plotted on the horizontal axis and the quantity supplied q is on vertical axis. Graph the
linear equation.
ii) Identify and interpret the meanings of slope p and q intercepts.
iii) Predict the weekly supply if the market price equals $35 per unit.
1
Expert's answer
2020-05-25T21:16:27-0400

Solution.

i)  If price p is plotted on the horizontal axis and the quantity supplied q is on vertical axis then the graph will be a line passing through the points with coordinates (15;30000) and (20;48000).



ii)The equation of the line has the form: "y=kx+b" were "k=tan\alpha."

Find "k" and "b" from the system of equations

"\begin{cases}
 30000=15k+b\\
 48000=20k+b
\end{cases};"

"\begin{cases}
 48000=20k+b \\
 30000=15k+b
\end{cases} -;"

"18000=5k;"

"k=3600;" "tan\alpha=3600;"

"30000=15\sdot3600+b \implies b=-24000;"

"q=3600p-24000 -" equation of this graph;

iii)If "p=35" $ then "q=3600\sdot35-24000=102000" units;

Answer:If p=35 $ then 102000 units.